OpenStudy (blasofia98):

If the amount of air in a spherical balloon is 288 cu. in., what is the radius of the balloon? a. 6 inches b. 7.3 inches c.72 inches

OpenStudy (mecharv):

The amount of air translates to the volume of the sphere... so what is the formula for volume of a sphere?

OpenStudy (blasofia98):

I don't know, I don't understand this question at all

OpenStudy (mecharv):

You dont know the volume of a sphere? :O

OpenStudy (mecharv):

I am shocked cause I forgot it too :( xD

OpenStudy (blasofia98):

This is a remedial class for me. Im just trying to get done with it as fast as i can so i can graduate

OpenStudy (mecharv):

I am just kidding, so yeah Volume(V) of a sphere is \[V=\frac{ 4 }{ 3 }Pir^3\]

OpenStudy (blasofia98):

okay

OpenStudy (mecharv):

V=288 So, \[288=\frac{ 4 }{ 3 }\times \pi \times r^3\] Pi=3.14 so just find r :)

OpenStudy (blasofia98):

Thank you!

OpenStudy (mecharv):

Anytime :)

OpenStudy (mecharv):

\[r^3=\frac{ 3 \times 288 }{ 4 \times 3.14 }\] \[r^3=\frac{ 864 }{ 12.56 }=68.789\] \[r=\sqrt[3]{789}=4.0973\]